# On the Feasibility of Simple Transformer for Dynamic Graph Modeling

### 25 Jan 2024 v1

@@ Abstract.
Dynamic graph modeling is crucial for understanding complex structures in web graphs, spanning applications in social networks, recommender systems, and more. Most existing methods primarily emphasize structural dependencies and their temporal changes. However, these approaches often overlook detailed temporal aspects or struggle with long-term dependencies. Furthermore, many solutions overly complicate the process by emphasizing intricate module designs to capture dynamic evolutions. In this work, we harness the strength of the Transformer’s self-attention mechanism, known for adeptly handling long-range dependencies in sequence modeling. Our approach offers a simple Transformer model tailored for dynamic graph modeling without complex modifications. We re-conceptualize dynamic graphs as a sequence modeling challenge and introduce an innovative temporal alignment technique. This technique not only captures the inherent temporal evolution patterns within dynamic graphs but also streamlines the modeling process of their evolution. As a result, our method becomes versatile, catering to an array of applications. Our model’s effectiveness is underscored through rigorous experiments on four real-world datasets from various sectors, solidifying its potential in dynamic graph modeling. The datasets and codes are available111https://anonymous.4open.science/r/SimpleDyG/.
@@ 1. Introduction
Graph-structured data are prevalent on the World Wide Web, such as social networks (Perozzi et al., 2014; Fan et al., 2019), recommender systems (Song et al., 2019; Wang et al., 2019b), article citation graphs (Xiong et al., 2017; Kanakia et al., 2019), dialogue systems (Lei et al., 2020; Liao et al., 2021), and so on. Thus, graph-based mining and learning have become fundamental tools in many Web applications, ranging from analyzing users’ behaviors ranging from the message-exchanging within social friendships, ratings/reviews on recommender platforms, publication/citation trends in the academic community, to multi-turn task-oriented dialogue. Traditionally, many works focus on static graphs characterized by fixed nodes and edges. However, many real-world graphs on the Web are intrinsically dynamic in nature, which continuously evolve over time (Sankar et al., 2020). That is, the nodes and their edges in such graphs are undergoing constant addition or reorganization based on some underlying patterns of evolution. For example, in a social network like UCI (Panzarasa et al., 2009), where nodes represent users and edges represent friend connections, users frequently exchange messages with their friends, and the social graph structure is constantly changing as new friendships are formed. To study this important class of graphs and their applications on the Web, we focus on dynamic graph modeling in this paper, aiming to capture the evolving patterns in a dynamic graph.
@@
Existing works for dynamic graph modeling mainly fall into two categories: discrete-time approaches (Sankar et al., 2020; Pareja et al., 2020) and continuous-time approaches (Trivedi et al., 2019; Xu et al., 2020; Wen and Fang, 2022; Cong et al., 2022) as shown in Figure 1(a) and 1(b), respectively. The former regards dynamic graphs as a sequence of snapshots over a discrete set of time steps. This kind of approach usually leverages structural modules such as graph neural networks (GNN) (Wu et al., 2020) to capture the topological information of graphs, and temporal modules such as recurrent neural networks (RNN) (Schuster and Paliwal, 1997) to learn the sequential evolution of dynamic graphs (Sankar et al., 2020). Meanwhile, the latter focuses on modeling continuous temporal patterns via specific temporal modules such as temporal random walk (Nguyen et al., 2018) or temporal kernel (De Vries and Principe, 1992), illustrated by Figure 1(b). Despite the achievements of previous works in dynamic graphs, there still exist some key limitations. First, the modeling of temporal dynamics on graphs is still coarse-grained or short-termed. On one hand, discrete-time approaches discard the fine-grained temporal information within the snapshot, which inevitably results in partial loss of temporal patterns. On the other hand, while continuous-time approaches retain full temporal details by mapping each interaction to a continuous temporal space, capturing long-term dependency within historical graph data still remains a difficult problem (Rossi et al., 2020; Yu et al., 2023). Second, the majority of the existing works rely extensively on the message-passing GNNs to encode the structural patterns in dynamic graphs. Although powerful in graph modeling, the message-passing mechanism shows inherent limitations such as over-smoothing (Chen et al., 2020) and over-squashing (Alon and Yahav, 2020) that become more pronounced as model depth increases, preventing deeper and more expressive architectures.
@@
In pursuit of addressing these limitations, we have witnessed the successful application of Transformer (Vaswani et al., 2017) and its variants in natural language processing (NLP) (Kenton and Toutanova, 2019; Brown et al., 2020) and computer vision (CV) (Dosovitskiy et al., 2020; Liu et al., 2021). The success is underpinned by two distinct advantages inherent to the Transformer architecture: as shown in Figure 1(c), it can naturally support a continuous sequence of data without the need for discrete snapshots, and its self-attention mechanism can capture long-term dependency (Vaswani et al., 2017), which are important factors for dynamic graph modeling. Transformer also presents a potentially better alternative to capturing topological information, as it is less or not affected by the over-smoothing and over-squashing issues associated with message-passing GNNs. Hence, in this work, we explore the feasibility of the Transformer architecture for dynamic graph modeling. In fact, we have observed a growing body of research trying to modify the Transformer for static graphs (Ying et al., 2021; Rampášek et al., 2022; Kim et al., 2022). Nonetheless, these studies primarily focus on integrating graph structural knowledge into the vanilla Transformer model, which generally still leverage message-passing GNNs as auxiliary modules to refine positional encoding and attention matrices based on graph-derived information (Min et al., 2022a). More recently, Ying et al. (2021) indicated that the pure Transformer architecture holds promise for graphs. However, all these previous Transformer-based approaches only focus on static graphs, leaving unanswered questions about the feasibility for dynamic graphs, as we elaborate below.
@@
The second challenge lies in the need to align temporal information across input sequences. Specifically, on dynamic graphs different input sequences converge on a common time domain—whether absolute points in time (e.g., 10am on 12 October 2023) or relative time intervals (e.g., a one-hour time window) convey the same across all sequences generated from different nodes’ history. In contrast, sequences for language modeling or static graphs lack such a universal time domain, and can be regarded as largely independent of each other. Thus, vanilla sequences without temporal alignment lack a way to differentiate variable time intervals and frequency information. For example, a bursty stream of interactions, happening over a short one-hour interval, has a distinct evolution pattern from a steady stream containing the same number of interactions, but happening over a period of one day.
Therefore, it becomes imperative to introduce a mechanism that infuses temporal alignment among distinct input sequences generated from the ego-graphs. To address this challenge, we carefully design special temporal tokens to align different input sequences in the time domain. The temporal tokens serve as indicators of distinct time steps that are globally recognized across all nodes and integrate them into the input sequences. While achieving the global alignment, local sequences from each node still retain the chronological order of the interactions in-between the temporal tokens, unlike traditional discrete-time approaches where temporal information within each snapshot is lost.
Based on the above insights, we propose a Simple Transformer architecture for Dynamic Graph modeling, named SimpleDyG. The word “simple” is a reference to the use of the original Transformer architecture without any modification, where the capability of dynamic graph modeling is simply and solely derived from constructing and modifying the input sequences. In summary, the contribution of our work is threefold.
We explore the potential of the Transformer architecture for modeling dynamic graphs. We propose a simple yet surprisingly effective Transformer-based approach for dynamic graphs, called SimpleDyG, without complex modifications.
We introduce a novel strategy to map a dynamic graph into a set of sequences, by considering the history of each node as a temporal ego-graph. Furthermore, we design special temporal tokens to achieve global temporal alignment across nodes, yet preserving the chronological order of interactions at a local level.
We conduct extensive experiments and analysis across four real-world Web graphs, spanning diverse applications domains on the Web. The empirical results demonstrate not only the feasibility, but also the superiority of SimpleDyG.
@@ 2.1. Dynamic Graph Learning
Current dynamic graph learning methods can be categorized into two primary paradigms: discrete-time approaches and continuous-time approaches. In discrete-time methods, dynamic graphs are treated as a series of static graph snapshots taken at regular time intervals. To model both structural and temporal aspects, these approaches integrate the GNNs with sequence models (RNNs or self-attention mechanisms) (Sankar et al., 2020; Pareja et al., 2020; Goyal et al., 2020; Taheri et al., 2019). For instance, DySAT Sankar et al. (2020) leverages Graph Attention Network (GAT) and self-attention as fundamental components for both structural and temporal modules. In contrast, EvolveGCN (Pareja et al., 2020) employed an RNN to evolve the Graph Convolutional Network (GCN) parameters. Nevertheless, they often fall short of capturing the granular temporal information. Consequently, the continuous-time approaches treat the dynamic graphs as sequences of interaction events at a specific timestamp. Some approaches model dynamic graph evolution as temporal random walks or causal anonymous walks (Nguyen et al., 2018; Wang et al., 2021). Another avenue of research focuses on time window encoding techniques integrated with graph structure modeling such as temporal graph attention used in TGAT (Xu et al., 2020) and TGN (Rossi et al., 2020) or MLP-Mixer layers applied in GraphMixer (Cong et al., 2022). Additionally, researchers also leverage temporal point processes treating the arrival of nodes/edges as discrete events (Trivedi et al., 2019; Ji et al., 2021; Wen and Fang, 2022). Despite the promise demonstrated by continuous-time approaches, it’s important to note that they come with limitations in capturing long-term dependencies originating from historical data.
@@
The differences between our work and the previous dynamic graph learning methods lie in two points. First, our method effectively mitigates long-term dependency challenges, leveraging the inherent advantages of the Transformer architecture. Second, our method preserves the chronological history of each ego node within the input sequences. The temporal alignment mechanisms among various ego networks empower our model to capture both global and local information within the dynamic graphs.
2.2. Transformers for Graphs
@@
Transformer architectures for graphs have emerged as a compelling alternative to conventional GNNs, aiming to mitigate issues like over-smoothing and over-squashing. Prior research focused on integrating graph information into the vanilla Transformer through diverse strategies. Some methods integrate GNNs as auxiliary components to bolster structural comprehension within the Transformer architecture (Rong et al., 2020; Kim et al., 2021). Others focus on enriching positional embeddings by spatial information derived from the graph. For instance, Graphormer (Ying et al., 2021) integrates the centrality, spatial and edge encoding into Transformer. Cai and Lam (2020) adopted distance embedding for tree-structured abstract meaning representation graph. Kreuzer et al. (2021) utilized the full Laplacian spectrum to learn the positional encoding for graph. There are also studies focus on refining attention mechanisms in Transformer for graph analysis. For instance, Min et al. (2022b) employed a graph masking attention mechanism to seamlessly inject graph-related priors into the Transformer architecture. Excepted for the complicated design, more recently, Kim et al. (2022) shed light on the effectiveness of pure Transformers in graph learning. Their approach treats all nodes and edges as independent tokens, severing as inputs for Transformer. Recently, Mao et al. (2023) proposed a Transformer based model for heterogeneous information networks. Node-level structure and heterogeneous relation are integrated into the attention mechanism.
@@
It’s worth noting that most of the previous works based on Transformers mainly focused on static graphs. Recently, Yu et al. (2023) introduced a Transformer based model designed for dynamic graph learning, which belongs to a contemporary work with ours. The difference lies in that they rely on complex designs for handling co-occurrence neighbors of different nodes and temporal interval encoding. In contrast, we explore the feasibility of a simple Transformer for dynamic graphs without the need for complex modifications.
@@ 3. PRELIMINARIES
In this section, we first illustrate the problem of dynamic graph modeling. Then we briefly introduce the main components of Transformer architecture.
@@ 3.1. Dynamic Graph Modeling
We define a dynamic graph as with a set of nodes , edges , a time domain and an input feature matrix . It can be characterized by a sequence of interacted links . Here, each tuple denotes a distinct interaction between nodes and at time , with representing the number of interactions within the temporal graph. Given the dynamic graph , we learn a model with parameter to capture the temporal evolution of the graph. The learned temporal representations can be used for different tasks such as node classification, link prediction and graph classification.
@@ 3.2. Transformer Architecture
We represent an input sequence as , where is the dimension of node features and is the hidden representation for token . The MHA module projects H to H sub-spaces denoted as:
where , , re the learnable parameter matrices, and their dimensions are set as .
The self-attention operation is performed using a scaled dot-product on the corresponding for each head:
where is learnable parameter matrix.
@@ 4. Proposed Approach
The overall framework of SimpleDyG is illustrated in Figure 2. Generally speaking, our framework is applied to a dynamic graph (Figure 2(a)), where multiple temporal links emerge at various time points. In order to capture the dynamic evolution, we begin by extracting temporal ego-graph for ego-node which contains the entire historical interactions as shown in Figure 2(b). These temporal graphs are subsequently transformed into sequences while preserving their chronological order. To incorporate temporal alignment among different ego-graphs, we segment the timeline into various time spans with the same temporal interval as in Figure 2(c). Then we add temporal tokens into the ego-sequence to make our model identify different time spans. Finally, these sequences are fed into a Transformer architecture to facilitate various downstream tasks.

### 27 Feb 2024 v2

@@ Abstract.
Dynamic graph modeling is crucial for understanding complex structures in web graphs, spanning applications in social networks, recommender systems, and more. Most existing methods primarily emphasize structural dependencies and their temporal changes. However, these approaches often overlook detailed temporal aspects or struggle with long-term dependencies. Furthermore, many solutions overly complicate the process by emphasizing intricate module designs to capture dynamic evolutions. In this work, we harness the strength of the Transformer’s self-attention mechanism, known for adeptly handling long-range dependencies in sequence modeling. Our approach offers a simple Transformer model, called SimpleDyG, tailored for dynamic graph modeling without complex modifications. We re-conceptualize dynamic graphs as a sequence modeling challenge and introduce a novel temporal alignment technique. This technique not only captures the inherent temporal evolution patterns within dynamic graphs but also streamlines the modeling process of their evolution. To evaluate the efficacy of SimpleDyG, we conduct extensive experiments on four real-world datasets from various domains. The results demonstrate the competitive performance of SimpleDyG in comparison to a series of state-of-the-art approaches despite its simple design.
@@ 1. Introduction
Graph-structured data are prevalent on the World Wide Web (Yu et al., 2024), such as social networks (Perozzi et al., 2014; Fan et al., 2019), recommender systems (Wu et al., 2020a, 2022b), citation graphs (Kanakia et al., 2019; Yu et al., 2023a, c), dialogue systems (Wu et al., 2022a; Liao et al., 2023), and so on. Thus, graph-based mining and learning have become fundamental tools in many Web applications, ranging from analyzing communication patterns within social friendships, to predicting users’ behavior on digital platforms and investigating citation trends in the academic community. Traditionally, many works focus on static graphs characterized by fixed nodes and edges. However, many real-world graphs on the Web are intrinsically dynamic, which continuously evolve over time (Sankar et al., 2020). That is, the nodes and their edges in such graphs are undergoing constant addition or reorganization based on some underlying patterns of evolution. For example, in a social network like UCI (Panzarasa et al., 2009), where nodes represent users and edges represent messages exchanged between users, new edges constantly emerge as users frequently exchange messages with their friends. To study this important class of graphs and their applications on the Web, we focus on dynamic graph modeling in this paper, aiming to capture the evolving patterns in a dynamic graph.
@@
Existing works for dynamic graph modeling mainly fall into two categories: discrete-time approaches (Sankar et al., 2020; Pareja et al., 2020) and continuous-time approaches (Trivedi et al., 2019; Xu et al., 2020; Wen and Fang, 2022; Cong et al., 2022), as shown in Figures 1(a) and 1(b), respectively. The former regards dynamic graphs as a sequence of snapshots over a discrete set of time steps. This kind of approach usually leverages structural modules such as graph neural networks (GNN) (Wu et al., 2020b) to capture the topological information of graphs, and temporal modules such as recurrent neural networks (RNN) (Schuster and Paliwal, 1997) to learn the sequential evolution of the snapshots (Sankar et al., 2020). Meanwhile, the latter focuses on modeling continuous temporal patterns via specific temporal modules such as temporal random walk (Nguyen et al., 2018) or temporal kernel (De Vries and Principe, 1992), illustrated by Figure 1(b). Despite the significant progress made in dynamic graph modeling, there still exist some key limitations. First, the modeling of temporal dynamics on graphs is still coarse-grained or short-termed. On one hand, discrete-time approaches discard the fine-grained temporal information within the snapshot, which inevitably results in partial loss of temporal patterns. On the other hand, while continuous-time approaches retain full temporal details by mapping each interaction to a continuous temporal space, capturing long-term dependency within historical graph data still remains a difficult problem (Rossi et al., 2020; Yu et al., 2023b). Second, the majority of the existing works rely extensively on the message-passing GNNs to encode the structural patterns in dynamic graphs. Although powerful in graph modeling, the message-passing mechanism shows inherent limitations such as over-smoothing (Chen et al., 2020) and over-squashing (Alon and Yahav, 2020) that become more pronounced as model depth increases, preventing deeper and more expressive architectures.
@@
In pursuit of addressing these limitations, we have been intrigued by the successful application of Transformer (Vaswani et al., 2017) and its variants in natural language processing (NLP) (Kenton and Toutanova, 2019; Brown et al., 2020; Liao et al., 2021a) and computer vision (CV) (Dosovitskiy et al., 2020; Liu et al., 2021). The success is underpinned by two distinct advantages inherent to the Transformer architecture: as shown in Figure 1(c), it can naturally support a continuous sequence of data without the need for discrete snapshots, and its self-attention mechanism can capture long-term dependency (Vaswani et al., 2017), which are important factors for dynamic graph modeling. Transformer also presents a potentially better alternative to capturing topological information, as it is less affected by the over-smoothing and over-squashing issues associated with message-passing GNNs. Hence, in this work, we explore the feasibility of the Transformer architecture for dynamic graph modeling. In fact, we have observed a growing body of research trying to modify the Transformer for static graphs (Ying et al., 2021; Rampášek et al., 2022; Kim et al., 2022). Nonetheless, these studies primarily focus on integrating graph structural knowledge into the vanilla Transformer model, which generally still leverage message-passing GNNs as auxiliary modules to refine positional encoding and attention matrices based on graph-derived information (Min et al., 2022a). More recently, Ying et al. (2021) have demonstrated that the pure Transformer architecture holds promise for graphs. However, all these previous Transformer-based approaches only focus on static graphs, leaving unanswered questions about the feasibility for dynamic graphs, as follows.
@@
The second challenge lies in the need to align temporal information across input sequences. Specifically, on dynamic graphs different input sequences converge within a common time domain—whether through absolute points in time (e.g., 10am on 12 October 2023) or relative time intervals (e.g., a one-hour window), with uniformity across all sequences generated from different nodes’ history. In contrast, sequences for natural language or static graphs lack such a universal time domain, and can be regarded as largely independent of each other. Thus, vanilla sequences without temporal alignment lack a way to differentiate variable time intervals and frequency information. For example, a bursty stream of interactions, happening over a short one-hour interval, has a distinct evolution pattern from a steady stream containing the same number of interactions, but happening over a period of one day.
Therefore, it becomes imperative to introduce a mechanism that infuses temporal alignment among distinct input sequences generated from the ego-graphs. To address this challenge, we carefully design special temporal tokens to align different input sequences in the time domain. The temporal tokens serve as indicators of distinct time steps that are globally recognized across all nodes, thereby unifying different input sequences. While achieving the global alignment, local sequences from each node still retain the chronological order of the interactions in-between the temporal tokens, unlike traditional discrete-time approaches where temporal information within each snapshot is lost.
Based on the above insights, we propose a Simple Transformer architecture for Dynamic Graph modeling, named SimpleDyG. The word “simple” is a reference to the use of the original Transformer architecture without any modification, where the capability of dynamic graph modeling is simply and solely derived from constructing and modifying the input sequences. In summary, the contribution of our work is threefold. (1) We explore the potential of the Transformer architecture for modeling dynamic graphs. We propose a simple yet surprisingly effective Transformer-based approach for dynamic graphs, called SimpleDyG, without complex modifications. (2) We introduce a novel strategy to map a dynamic graph into a set of sequences to improve the scalability, by considering the history of each node as a temporal ego-graph. Furthermore, we design special temporal tokens to achieve global temporal alignment across nodes, yet preserving the chronological order of interactions at a local level. (3) We conduct extensive experiments and analyses across four real-world Web graphs, spanning diverse domains on the Web. The empirical results demonstrate not only the feasibility, but also the superiority of SimpleDyG.
@@ 2. Related Work
Dynamic Graph Learning. Current dynamic graph learning methods can be categorized into discrete-time and continuous-time approaches. In discrete-time methods, dynamic graphs are treated as a series of static graph snapshots taken at regular time intervals. To model both structural and temporal aspects, these approaches integrate the GNNs with sequence models, such as RNNs or self-attention mechanisms (Sankar et al., 2020; Pareja et al., 2020; Goyal et al., 2020; Taheri et al., 2019). For instance, DySAT (Sankar et al., 2020) leverages a graph attention network and self-attention as fundamental components for both structural and temporal modules, whereas EvolveGCN (Pareja et al., 2020) employs an RNN to evolve the parameters of graph convolutional networks. Nevertheless, they often fall short of capturing the granular temporal information. In contrast, the continuous-time approaches consider every interaction event at each specific timestamp. Some approaches model dynamic graph evolution as temporal random walks or causal anonymous walks (Nguyen et al., 2018; Wang et al., 2021). Another line of research focuses on time encoding techniques, which integrate with graph structure modeling, such as the temporal graph attention used in TGAT (Xu et al., 2020) and TGN (Rossi et al., 2020), or MLP-Mixer layers applied in GraphMixer (Cong et al., 2022). Additionally, researchers also leverage temporal point processes to capture the graph formation process (Trivedi et al., 2019; Ji et al., 2021; Wen and Fang, 2022). Despite the promise demonstrated by continuous-time approaches, it is important to note that they come with limitations in capturing long-term dependencies within historical data.
@@
The differences between our work and the previous dynamic graph learning methods lie in two points. First, our method effectively mitigates the long-term dependency challenge, leveraging the inherent advantages of the Transformer architecture. Second, our method preserves the chronological history of each input sequence. In particular, the temporal alignment mechanism synchronizes different input sequences, empowering our model to capture both global and local information within the dynamic graphs.
@@
Transformers for Graphs. Transformer architectures for graphs have emerged as a compelling alternative to conventional GNNs, aiming to mitigate issues like over-smoothing and over-squashing. Prior research has focused on integrating graph information into the vanilla Transformer through diverse strategies. Some methods integrate GNNs as auxiliary components to bolster structural comprehension within the Transformer architecture (Rong et al., 2020; Kim et al., 2021). Others focus on enriching positional embeddings by spatial information derived from the graph. For instance, Graphormer (Ying et al., 2021) integrates the centrality, spatial and edge encoding into Transformers, whereas Cai and Lam (2020) adopt distance embedding for tree-structured abstract meaning representation graph, and Kreuzer et al. (2021) utilize the full Laplacian spectrum to learn the positional encoding. There are also studies on refining the attention mechanism in Transformers for graph modeling. For instance, Min et al. (2022b) employ a graph masking attention mechanism to seamlessly inject graph-related priors into the Transformer architecture. More recently, Kim et al. (2022) have shed light on the effectiveness of pure Transformers in graph learning without complex designs. Their approach treats all nodes and edges as independent tokens, serving as inputs for the Transformer. Besides, Mao et al. (2023) propose a Transformer-based model for heterogeneous graphs, integrating local structures and heterogeneous relations into the attention mechanism.
@@
It is worth noting that most of the previous works based on Transformers mainly deal with static graphs. Recently, Yu et al. (2023b) have introduced a Transformer-based model designed for dynamic graph learning, which is contemporary with our work. The difference lies in that they rely on complex designs for capturing co-occurrence neighbors of different nodes and encoding temporal intervals. In contrast, we explore a simple Transformer for dynamic graphs without the need for complex modifications.
@@ 3. PRELIMINARIES
We first define the problem of dynamic graph modeling. Then, we briefly introduce the background of the Transformer architecture.
@@ 3.1. Dynamic Graph Modeling
We define a dynamic graph as with a set of nodes , edges , a time domain and an input feature matrix . It can be characterized by a sequence of time-stamped edges . Here, each tuple denotes a distinct interaction between nodes and at time , with representing the total number of interactions in the dynamic graph. Given the dynamic graph , we learn a model with parameter to capture the temporal evolution of the graph. The temporal representations encoded by the learned model can be used for different tasks such as node classification, link prediction and graph classification.
@@ 3.2. Transformer Architecture
We represent an input sequence as , where is the hidden representation for token , and is the dimension of the representations. The MHA module projects H to a triplet , as follows.
where , , are learnable weights, with . Overall, such projections are performed, resulting in for . The self-attention operation is then applied to each triplet:
where is a learnable weight matrix.
@@ 4. Proposed Approach
The overall framework of SimpleDyG is illustrated in Figure 2. Our framework is applied to a dynamic graph (Figure 2(a)), where multiple temporal links emerge at various time points. In order to capture the dynamic evolution, we begin by extracting a temporal ego-graph for each ego-node, which contains the entire historical interactions as shown in Figure 2(b). These temporal graphs are subsequently transformed into sequences while preserving their chronological order. To incorporate temporal alignment among different ego-graphs, we segment the timeline into various time spans with the same temporal interval as in Figure 2(c). Then, we add temporal tokens into the ego-sequence to identify different time spans. Finally, these sequences are fed into a Transformer architecture to facilitate various downstream tasks.